Fermi National Laboratory


Accelerator Update

Monday August 9
The day shift began with Tevatron (Tev) store #3717 colliding, stacking, and with beam to MiniBoone and Switchyard 120.

At 5:26 PM the Antiproton Source (Pbar) stopped stacking and went to reverse protons for studies, this lasted ~ 5 hours.

Tuesday August 10
The Midnight shift began with Store #3717 colliding, stacking, and with beam to MiniBoone, MCenter, and MTest.

At 6:06 AM the Main Injector (MI) power supply loop dropped out. The problem was traced to MI20 and experts were called to investigate.

At 10:33 AM the Pbar stack was lost when QDF went to 0 current.

After repairs to the MI-20 power supply, the MI was back on at 10:42 AM.

Store #3717 was terminated at 12:10 PM.

Repairs to QDF were completed by 2:58 PM and Pbar returned to stacking at 3:43 PM.

CDF made a controlled access at 4:00 PM, this lasted until 5:00 PM.

The Tevatron Department performed Tev studies all evening while waiting for Pbar to build a stack for the next shot setup.

Wednesday August 11
The Midnight shift began with Tev still in a study mode, stacking, and with beam to MiniBoone and Switchyard 120.

The Tev went into mixed mode shot setup at 4:15 AM, with store #3723 colliding at 6:59 AM with an initial average luminosity of 61.77 E30.

At 5:35 the Meson Low Conductivity Water (LCW) system at Service building MS2 tripped off causing power supplies in Meson to trip.

The MS2LCW system was returned to service at 7:10 AM after replacing an air compressor for the system.

Today's Plans
Continue to stack Pbars and maintain Tev store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/9/2004   email Fermilab